# Intuitionistic Fuzzy Universum Twin Support Vector Machine for Imbalanced Data

### v1

@@ 2.1 Intuitionistic Fuzzy Membership Scheme
2.2 Universum Twin Support Vector Machine (UTSVM)
$\underset{w_{1},b_{1},\xi_{1},\psi_{1}}{min}\hskip 14.22636pt~{}$ $\frac{1}{2}\|X_{1}w_{1}+e_{1}b_{1}\|^{2}+c_{1}e_{2}^{t}\xi_{1}+c_{u}e_{u}^{t}\psi_{1}$
s.t. $-\left(X_{2}w_{1}+e_{2}b_{1}\right)\geq e_{2}-\xi_{1},$
$\left(Uw_{1}+e_{u}b_{1}\right)\geq(-1+\epsilon)e_{u}-\psi_{1},$
$\hskip 17.07182pt\xi_{1}\geq 0,\psi_{1}\geq 0,$ (7)
and
$\underset{w_{2},b_{2},\xi_{2},\psi_{2}}{min}\hskip 14.22636pt~{}$ $\frac{1}{2}\|X_{2}w_{2}+e_{2}b_{2}\|^{2}+c_{2}e_{1}^{t}\xi_{2}+c_{u}e_{u}^{t}\psi_{2}$
s.t. $\left(X_{1}w_{2}+e_{1}b_{2}\right)\geq e_{1}-\xi_{2},$
$-\left(Uw_{2}+e_{u}b_{2}\right)\geq(-1+\epsilon)e_{u}-\psi_{2},$
$\hskip 17.07182pt\xi_{2}\geq 0,\psi_{2}\geq 0,$ (8)
here, $c_{i},(i=1,2,u)$ are positive parameters, $\xi_{1},\xi_{2},\psi_{1},\psi_{2}$ are slacks variables. The dual formulations of problems (2.2) and (2.2) can be obtaind as:
$\underset{\alpha_{1},\mu_{1}}{max}\hskip 14.22636pt~{}$ $e_{2}^{t}\alpha_{1}-\frac{1}{2}(\alpha_{1}^{t}E-\mu_{1}^{t}F)(GG^{t})^{-1}(E^{t}\alpha_{1}-F^{t}\mu_{1})+(\epsilon-1)e_{u}^{t}\mu_{1}$
s.t $0\leq\mu_{1}\leq c_{u},$
$0\leq\alpha_{1}\leq c_{1},$
and
$\underset{\alpha_{2},\mu_{2}}{max}\hskip 14.22636pt~{}$ $e_{1}^{t}\alpha_{2}-\frac{1}{2}(\alpha_{2}^{t}G-\mu_{2}^{t}F)(EE^{t})^{-1}(G^{t}\alpha_{2}-F^{t}\mu_{2})+(\epsilon-1)e_{u}^{t}\mu_{2}$
s.t $0\leq\mu_{2}\leq c_{u},$
$0\leq\alpha_{2}\leq c_{2},$
here, $G=[X_{1},e_{1}]$, $E=[X_{2},e_{2}]$, $F=[U,e_{u}]$ and $\alpha_{1}$, $\alpha_{2}$, $\mu_{1}$, $\mu_{2}$ are Lagrange multipliers of appropriate dimensions. The solution of (2.2) and (2.2) yields the separating hyperplane,
[…]
3 Proposed Intuitionistic Fuzzy Universum Twin Support Vector Machine for Imbalanced Data (IFUTSVM-ID)
In this section, we provide a detailed mathematical formulation of the proposed IFUTSVM-ID model tailored for linear and nonlinear cases. In case of class imbalance, the majority class holds a significant amount of information regarding the data distribution. When dealing with a majority-class hyperplane, we increase the number of data points in the universum data to proper information about the data. In contrast to UTSVM, which does not address the issue of class imbalance and considers the universum as not belonging to either of the two classes. Constraints on the universum samples are included such that their sizes are equivalent to the difference between the number of samples in the classes. To construct the hyperplane for the majority class, a constraint is applied to ensure that the distance of universum data points from the majority (negative) class must be $(1-\epsilon)$. In addition to giving prior information of data, the majority class hyperplane is not positioned too close to the data point of the minority (positive) class. Consequently, this scheme reduces the inclination of the classifier towards the majority class, leading to improved generalization performance.
3.1 Linear IFUTSVM-ID
$\underset{w_{1},b_{1},\xi_{1},\lambda_{1}}{min}\hskip 14.22636pt~{}$ $\frac{1}{2}\|X_{1}w_{1}+e_{1}b_{1}\|^{2}+\frac{c_{3}}{2}(\|w_{1}\|^{2}+b_{1}^{2})+c_{1}s_{2}^{t}\xi_{1}+c_{u}e_{g}^{t}\lambda_{1}$
s.t. $-\left(X_{2}^{*}w_{1}+e_{2}b_{1}\right)\geq e_{2}-\xi_{1},$
$\left(U^{*}w_{1}+e_{g}b_{1}\right)\geq(-1+\epsilon)e_{g}-\lambda_{1},$
$\hskip 17.07182pt\xi_{1}\geq 0,\lambda_{1}\geq 0,$ (18)
and
$\underset{w_{2},b_{2},\xi_{2},\lambda_{2}}{min}\hskip 14.22636pt~{}$ $\frac{1}{2}\|X_{2}w_{2}+e_{2}b_{2}\|^{2}+\frac{c_{4}}{2}(\|w_{2}\|^{2}+b_{2}^{2})+c_{2}s_{1}^{t}\xi_{2}+c_{u}e_{d}^{t}\lambda_{2}$
s.t. $\left(X_{1}w_{2}+e_{1}b_{2}\right)\geq e_{1}-\xi_{2},$
$\left(Uw_{2}+e_{d}b_{2}\right)\geq(\epsilon-1)e_{d}-\lambda_{2},$
$\hskip 17.07182pt\xi_{2}\geq 0,\lambda_{2}\geq 0,$ (19)
The Lagrangian of equation (3.1) is written as:
$L(\Theta)=~{}$ $\frac{1}{2}\|X_{1}w_{1}+e_{1}b_{1}\|^{2}+\frac{c_{3}}{2}(\|w_{1}\|^{2}+b_{1}^{2})+c_{1}s_{2}^{t}\xi_{1}+c_{u}e_{g}^{t}\lambda_{1}$
$-\alpha^{t}(-\left(X_{2}^{*}w_{1}+e_{2}b_{1}\right)-e_{2}+\xi_{1})$
By employing the K.K.T. conditions to equation (3.1), we obtain
$X_{1}^{t}(X_{1}w_{1}+e_{1}b_{1})+c_{3}w_{1}+X_{2}^{*t}\alpha-U^{*t}\beta=0,$ (21)
$e_{1}^{t}(X_{1}w_{1}+e_{1}b_{1})+c_{3}b_{1}+e_{2}^{t}\alpha-e_{g}^{t}\beta=0,$ (22)
$c_{1}s_{2}-\alpha-\mu=0,$ (23)
$c_{u}e_{g}-\beta-\psi=0,$ (24)
$\alpha^{t}(-\left(X_{2}^{*}w_{1}+e_{2}b_{1}\right)-e_{2}+\xi_{1})=0,$ (25)
$\mu^{t}\xi_{1}=0,$ (27)
$\gamma^{t}\lambda_{1}=0.$ (28)
Since $\mu>0$ and $\gamma>0,$ (23) and (24) reduce to,
$0\leq\alpha\leq c_{1}s_{2},$
$0\leq\beta\leq c_{u}e_{g}.$
Rewriting (21) and (22), we obtain
$\left(E^{t}E+c_{3}I\right)\binom{w_{1}}{b_{1}}+F^{*t}\alpha-G^{*t}\beta$
[…]
With the help of equation (3.1) and the K.K.T conditions (21) - (28), the dual form of equation (3.1) can be obtained as follows:
$\underset{\alpha,\beta}{max}\hskip 14.22636pt~{}$ $-\frac{1}{2}\left(\alpha^{t}F^{*}-\beta^{t}G^{*}\right)(E^{t}E+c_{3}I)^{-1}(F^{*t}\alpha-G^{*t}\beta)+(\epsilon-1)e_{g}^{t}\beta+e_{1}^{t}\alpha$
s.t. $0\leq\alpha\leq c_{1}s_{2},$
$0\leq\beta\leq c_{u}e_{g},$ (31)
and
$\underset{\eta,\theta}{max}\hskip 14.22636pt~{}$ $-\frac{1}{2}\left(\eta^{t}E+\theta^{t}G\right)(F^{t}F+c_{4}I)^{-1}(E^{t}\eta+G^{t}\theta)+(1-\epsilon)e_{g}^{t}\theta+e_{1}^{t}\eta$
s.t. $0\leq\eta\leq c_{2}s_{1},$
$0\leq\theta\leq c_{u}e_{d},$ (32)
After solving (3.1) and (3.1), we can acquire the separating hyperplanes as follows
[…]
3.2 Non-linear IFUTSVM-ID
$\underset{w_{1},b_{1},\xi_{1},\lambda_{1}}{min}\hskip 14.22636pt~{}$ $\frac{1}{2}\|\mathscr{K}(X_{1},D^{t})w_{1}+e_{1}b_{1}\|^{2}+\frac{c_{3}}{2}(\|w_{1}^{2}\|+b_{1}^{2})+c_{1}s_{2}^{t}\xi_{1}+c_{u}e_{g}^{t}\lambda_{1}$
$\hskip 17.07182pt\xi_{1}\geq 0,\lambda_{1}\geq 0,$ (36)
and
$\underset{w_{2},b_{2},\xi_{2},\lambda_{2}}{min}\hskip 14.22636pt~{}$ $\frac{1}{2}\|\mathscr{K}(X_{2},D^{t})w_{2}+e_{2}b_{2}\|^{2}+\frac{c_{4}}{2}(\|w_{2}^{2}\|+b_{2}^{2})+c_{2}s_{1}^{t}\xi_{2}+c_{u}e_{d}^{t}\lambda_{2}$
s.t. $\left(\mathscr{K}(X_{1},D^{t})w_{2}+e_{1}b_{2}\right)\geq e_{1}-\xi_{2},$
$\hskip 17.07182pt\xi_{2}\geq 0,\lambda_{2}\geq 0,$ (37)
The Wolfe duals of the problems of (3.2) and (3.2) can be obtained as:
$\underset{\alpha,\beta}{max}\hskip 14.22636pt~{}$ $-\frac{1}{2}\left(\alpha^{t}F^{*}-\beta^{t}G^{*}\right)(E^{t}E+c_{3}I)^{-1}(F^{*t}\alpha-G^{*t}\beta)+(\epsilon-1)e_{g}^{t}\beta+e_{1}^{t}\alpha$
s.t. $0\leq\alpha\leq c_{1}s_{2},$
$0\leq\beta\leq c_{u}e_{g},$ (38)
and
$\underset{\eta,\theta}{max}\hskip 14.22636pt~{}$ $-\frac{1}{2}\left(\eta^{t}E+\theta^{t}G\right)(F^{t}F+c_{4}I)^{-1}(E^{t}\eta+G^{t}\theta)+(1-\epsilon)e_{g}^{t}\theta+e_{1}^{t}\eta$
s.t. $0\leq\eta\leq c_{2}s_{1},$
$0\leq\theta\leq c_{u}e_{d},$ (39)
After solving (3.2) and (3.2), the separating hyperplanes can be attained as,
[…]
4 Computational Complexity
@@
3:  To obtain the optimal hyperplanes for each class, solve equations (3.1) and (34) for the linear case, or equations (40) and (41) for the nonlinear case.
4:  Testing sample is classified into class $+1$ or $-1$ using (35) or (42) for linear and non-linear case, respectively.
5 Experimental results
To test the efficiency of proposed IFUTSVM-ID model, we compare their performance against baseline models, including UTSVM [24], RUTSVM-CIL [33], IF-USVM [15], IFW-LSTSVM [37], and CGFTSVM-ID [16]. We utilize the publicly available KEEL [7] benchmark datasets under different noise levels. Furthermore, to assess the effectiveness of the proposed IFUTSVM-ID model, we utilized it for the diagnosis of Alzheimer’s Disease (AD). AD dataset is accessible through the Alzheimer’s Disease Neuroimaging Initiative (ADNI) (adni.loni.usc.edu).
5.1 Experimental Setup
@@
Table 1: Comparison of the performance of the proposed IFUTSVM-ID model with the baseline models based on classification ACC on KEEL datasets.
@@
5.2 Evaluation on KEEL datasets
In this subsection, we present an intricate analysis involving a comparison of the proposed IFUTSVM-ID model against baseline UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models. This comparison is carried out across $46$ KEEL benchmark datasets spanning diverse domains. Table 1 presents the classification ACC, along with the optimal parameters of the proposed IFUTSVM-ID model compared to the existing models. The average ACC of the existing UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models are $83.52\%$, $81.15\%$, $80.61\%$, $79.44\%$, and $86.19\%$ respectively, whereas the average ACC of the proposed IFUTSVM-ID model is $87.70\%$. The proposed IFUTSVM-ID secured the top position in terms of average ACC. This observation strongly emphasizes the significant superiority of the proposed models over baseline models. Depending solely on average ACC as a single metric could pose problems, as outstanding performance on specific datasets may obscure insufficient performance on others. To address this concern, it becomes crucial to rank each model individually with respect to each dataset, enabling a thorough evaluation of their respective capabilities. In the ranking scheme [6], the model with the lowest performance on a dataset receives a higher rank, whereas the model achieving the best performance is assigned a lower rank. In the evaluation of $p$ models across $N$ datasets, the rank of the $j^{th}$ model on the $i^{th}$ dataset can be denoted as $R_{j}^{i}$. Then, $R_{j}=\frac{1}{N}\sum_{i=1}^{N}R_{j}^{i}$ determined the average rank of the model. The average rank of the proposed IFUTSVM-ID with the existing UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models are $3.71$, $4.24$, $4.47$, $4.33$, $2.76$ and $1.54$, respectively. The proposed IFUTSVM-ID model attained the lowest average rank among all the models. Since a lower rank indicates a better-performing model, the proposed IFUTSVM-ID emerged as the top-performing model. We now proceed to conduct statistical tests to determine the significance of the results. Firstly, we employ the Friedman test [6] to determine whether there are significant differences among the models. Under the null hypothesis, it is assumed that the average rank of the models is equivalent, indicating equal performance. The Friedman statistic is given by: $\chi_{F}^{2}=\frac{12N}{p(p+1)}\left[\sum_{i=1}R_{j}^{2}-\frac{p(p+1)^{2}}{4}\right]$. However, the Friedman statistic tends to be overly cautious. To address this, the better statistic [12] is given as follows: $F_{F}=\frac{(N-1)\chi_{F}^{2}}{N(p-1)-\chi_{F}^{2}}$, where $F_{F}$ follows an $F$-distribution with d.o.f $((p-1),(p-1)(N-1))$. For $p=6$ and $N=46$, we get $\chi_{F}^{2}=91.48$ and $F_{F}=29.72$. From the $F$-distribution table $F_{F}(5,225)=2.2542$ at a $5\%$ level of significance. The null hypothesis rejected as $F_{F}>2.2542$. Therefore, significant disparities exist among the models. Subsequently, we utilize the Nemenyi post hoc test [6] to scrutinize the pairwise distinctions between the models. The critical difference $(C.D.)$ is computed at a significance level of $5\%$ as $C.D.=q_{\alpha}\sqrt{\frac{p(p+1)}{6N}}$, where $q_{\alpha}$ denotes the critical value obtained from the distribution table for the two-tailed Nemenyi test. Referring to the statistical $F$-distribution table, at a significance level of $5\%$, we find $q_{\alpha}=2.850$, which leads to a critical difference of $C.D.=1.11$. The average differences in rank between the proposed IFUTSVM-ID model and the baseline UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models are $2.17$, $2.70$, $2.93$, $2.79$, and $2.76$, respectively. The Nemenyi post hoc test reveals significant differences between the proposed IFUTSVM-ID model and the baseline models. Considering all these findings, we infer that the proposed IFUTSVM-ID model demonstrates superior performance when compared to the baseline models.
Table 2: Win-tie-loss of proposed IFUTSVM-ID model along with baseline models on KEEL datasets.
UTSVM [24] RUTSVM-CIL [33] IF-USVM [15] IFW-LSTSVM [37] CGFTSVM-ID [16]
RUTSVM-CIL [33] [15, 5, 26]
IF-USVM [15] [14, 7, 25] [18, 3, 25]
IFW-LSTSVM [37] [18, 3, 25] [21, 1, 24] [23, 3, 20]
CGFTSVM-ID [16] [30, 3, 13] [37, 3,6] [37, 1, 8] [32, 5, 9]
IFUTSVM-ID [36, 5, 5] [38, 6,2] [41, 3, 2] [43, 2, 1] [36, 6, 4]
Furthermore, to analyze the models, we use a pairwise win-tie-loss (W-T-L) sign test. In W-T-L, we determine the number of datasets where each classifier achieves a win, tie, or loss. In the test, if the null hypothesis is true, it means that the two models have the same level of performance. Therefore, each model is expected to perform well on exactly half of the total number of datasets, denoted as $N/2$. If one model has won on a number of datasets that is equal to or greater than $N/2+1.96\sqrt{N}/2$, it is considered significantly different from the other model. If there is a tie in the comparison between the two models, the number is divided equally between them. Table 2 shows the outcomes of the pairwise analysis of the models, which is obtained through simple calculations. With a sample size of $N=46$, the two models are considered statistically significant and distinct if one of the models wins on atleast $29.65$ datasets. Table 2 shows the discernible discrepancies between the classifiers, determined through the W-T-L is conducted on KEEL datasets. Table 2 clearly indicates that the proposed IFUTSVM-ID model exhibits significant superiority compared to the baseline model.
5.3 Evaluation on KEEL Datasets with Added Label Noise
The evaluation conducted using KEEL datasets reflects real-world scenarios. However, it’s crucial to acknowledge that data impurities or noise may stem from various factors. To showcase the effectiveness of the proposed IFUTSVM-ID model, especially under challenging conditions, we deliberately introduced label noise to certain datasets. We selected five datasets, namely ecoli-0-1_vs_5, aus, abalone9-18, checkerboard_Data, and brwisconsin to assess the robustness of the models. To ensure fairness in model evaluation, we intentionally selected two datasets where the proposed IFUTSVM-ID model did not achieve the highest performance and three datasets where they achieved comparable results to an existing model with varying levels of label noise. For a comprehensive analysis, we introduced label noise at various levels, including $5\%$, $10\%$, $15\%$, and $20\%$, intentionally corrupting the labels of these datasets. Table 3 displays the accuracies of all models for the selected datasets with $5\%$, $10\%$, $15\%$, and $20\%$ noise. Consistently, the proposed IFUTSVM-ID model demonstrates superior performance over baseline models, exhibiting higher ACC. Significantly, they maintain this leading performance despite the presence of noise. The average ACC of the proposed IFUTSVM-ID on the abalone9-18 dataset at various noise levels is $76.83\%$, surpassing the performance of the baseline models. On the checkerboard_Data dataset and ecoli-0-1_vs_5, the average ACC of the proposed model at different noise levels is $84.18$ and $78.47\%$, respectively. On the aus dataset, the proposed model achieves an average ACC of $85.05\%$, which is slightly lower than the average ACC of the IF-USVM model. However, the ACC of the proposed IFUTSVM-ID model reaches $86.47\%$ at a $0\%$ level of label noise. At each noise level, IFUTSVM-ID emerges as the top performer, with an overall average ACC of $79.79\%$. By subjecting the model to rigorous conditions, we aim to showcase the exceptional performance and superiority of the proposed IFUTSVM-ID model, especially in challenging scenarios. The above findings underscore the significance of the proposed IFUTSVM-ID model as resilient solutions, capable of performing well in demanding conditions characterized by noise and impurities.
Table 3: Performance comparison of the proposed IFUTSVM-ID with the baseline models on KEEL datasets with label noise.
$ACC(\%)$ $ACC(\%)$ $ACC(\%)$ $ACC(\%)$ $ACC(\%)$ $ACC(\%)$
abalone9-18 $5\%$ $76.8$ $64.84$ $75.39$ $75.89$ $75.89$ $77.56$
$10\%$ $72.69$ $64.84$ $59.36$ $74.89$ $75.26$ $75.97$
$15\%$ $79.5$ $70.32$ $75.39$ $75.89$ $75.43$ $73.01$
$20\%$ $76.35$ $77.17$ $79$ $79.37$ $79.04$ $80.78$
Average ACC $76.34$ $69.29$ $72.29$ $76.51$ $76.41$ $76.83$
aus $5\%$ $84.37$ $85.51$ $85.99$ $85.51$ $85.47$ $86.98$
$10\%$ $85.41$ $85.51$ $85.99$ $69.08$ $84.37$ $86.47$
$15\%$ $83.92$ $67.15$ $85.51$ $65.7$ $84.41$ $84.54$
$20\%$ $85.44$ $59.42$ $85.51$ $82.61$ $82.61$ $82.19$
Average ACC $84.79$ $74.4$ $85.75$ $75.73$ $84.22$ $85.05$
brwisconsin $5\%$ $77.55$ $65.69$ $78.04$ $77.55$ $77.55$ $75.69$
$10\%$ $76.57$ $65.69$ $77.06$ $75.1$ $76.57$ $74$
$15\%$ $72.65$ $79.71$ $75.59$ $74.12$ $72.16$ $73.63$
$20\%$ $74.12$ $76.57$ $76.57$ $75.1$ $74.61$ $74.31$
Average ACC $75.22$ $71.92$ $76.82$ $75.47$ $75.22$ $74.41$
checkerboard_Data $5\%$ $82.37$ $81.51$ $81.99$ $82.51$ $85.47$ $85.51$
$10\%$ $85.41$ $85.51$ $85.99$ $69.08$ $86.37$ $86.47$
$15\%$ $82.92$ $67.15$ $80.51$ $65.7$ $81.41$ $84.54$
$20\%$ $80.44$ $59.42$ $80.51$ $79.61$ $80.61$ $80.19$
Average ACC $82.79$ $73.4$ $82.25$ $74.23$ $83.47$ $84.18$
ecoli-0-1_vs_5 $5\%$ $73.06$ $79.17$ $78.89$ $73.06$ $75.83$ $79.17$
$10\%$ $74.44$ $79.17$ $75.99$ $73.06$ $75.83$ $77.78$
$15\%$ $73.06$ $78.89$ $68.06$ $73.06$ $74.44$ $72.22$
$20\%$ $88.89$ $87.22$ $87.22$ $83.06$ $84.44$ $84.72$
Average ACC $77.36$ $81.11$ $77.54$ $75.56$ $77.64$ $78.47$
Overall Average ACC $79.3$ $74.02$ $78.93$ $75.5$ $79.39$ $79.79$
5.4 Evaluation on ADNI dataset
Table 4: Performance comparison of the proposed IFUTSVM-ID along with the baseline models based on classification ACC on AD and BC datasets.
Average ACC $76.12$ $66.92$ $69.48$ $56$ $76.81$ $77.57$
The ACC values for the proposed IFUTSVM-ID model along with the existing models for AD diagnosis, are presented in Table 4. With an average ACC of $77.57\%$, the proposed IFUTSVM-ID emerges as the top classifier in our analysis. The average ACC of the baseline UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models are $76.12$, $66.92$, $69.48$, $56$, and $76.81$, respectively. Compared to the second-top model, CGFTSVM-ID, the proposed IFUTSVM-ID model exhibits an average ACC that is approximately $0.76\%$ higher. The proposed IFUTSVM-ID achieves the highest ACC of $89.87\%$ for the CN vs AD case. In the CN vs MCI case, the proposed IFUTSVM-ID emerges as the top performer, with average accuracies of $70.85\%$. Therefore, the proposed IFUTSVM-ID consistently demonstrates superior performance by achieving high ACC across various cases, establishing its prominence among the models. Furthermore, to illustrate the comparison between the proposed IFUTSVM-ID model and the baseline UTSVM, RUTSVM-CIL, IF-USVM, IFW-LSTSVM, and CGFTSVM-ID models in terms of specificity, sensitivity, and precision, we plotted the bar graphs shown in Figure 2. The performance of the proposed IFUTSVM-ID model is almost comparable to that of the baseline models. The overall findings emphasize the effectiveness of the proposed models in discerning various cognitive states.
6 Conclusion
In this paper, we proposed the intuitionistic fuzzy universum twin support vector machine for imbalanced data (IFUTSVM-ID) to address imbalanced datasets in the presence of noise and outliers. Unlike UTSVM, which employs a uniform weighting approach to generate the optimal classifier, the proposed IFUTSVM-ID model employs an intuitionistic fuzzy membership scheme to compute the final parameters. As a result, the proposed IFUTSVM-ID achieves superior generalization performance by incorporating prior information into the training process. The approach integrating undersampling with oversampling using universum data has been found to be beneficial in classifying imbalanced datasets. One of the key advantages of incorporating a regularization term into the primal optimization formulation of IFUTSVM-ID is its adherence to the SRM principle. This approach bypasses the need for matrix inversions, thereby streamlining computational processes. Furthermore, it effectively mitigates overfitting concerns, enhancing the model’s generalization capabilities. We evaluate the proposed IFUTSVM-ID model on $46$ KEEL datasets to experimentally validate the performance. Statistical evaluations, including metrics such as accuracy (ACC), ranking scheme analysis, Win-tie-loss sign test, Friedman test, and Nemenyi post hoc test, conclusively demonstrate that the proposed IFUTSVM-ID model exhibits superior performance compared to other classifiers. Also, we employed the proposed IFUTSVM-ID model for the diagnosis of Alzheimer’s disease datasets. Upon analysis, it has been ascertained that the proposed IFUTSVM-ID model demonstrates the highest level of classification performance. In future, there is a possibility of examining the selection of universum data as it has an impact on the performance of the model. Further, to enhance the performance of our IFUTSVM-ID model, various undersampling and oversampling techniques can be employed.
Acknowledgment

### v2

@@ Acknowledgment
S.II Algorithm and Computational Complexity Analysis of the proposed IFUTSVM-ID model
@@
3:To obtain the optimal hyperplanes for each class, solve equations (26) and (30) for the linear case, or equations (36) and (37) for the nonlinear case of the main file.
4:Testing sample is classified into class $+1$ or $-1$ using (31) or (38) for linear and non-linear case of the main file, respectively.
S.III Experimental Setup and Experimental Results
In this section, we give a detailed description of the experimental setup and the experimental results on KEEL datasets with label noise for each model including baseline and proposed.
S.III-A Experimental Setup
@@
S.III-B Evaluation on KEEL Datasets with Added Label Noise
The evaluation conducted using KEEL datasets reflects real-world scenarios. However, it’s crucial to acknowledge that data impurities or noise may stem from various factors. To showcase the effectiveness of the proposed IFUTSVM-ID model, especially under challenging conditions, we deliberately introduced label noise to certain datasets. We selected five datasets, namely ecoli-0-1_vs_5, aus, abalone9-18, checkerboard_Data, and brwisconsin to assess the robustness of the models. To ensure fairness in model evaluation, we intentionally selected two datasets where the proposed IFUTSVM-ID model did not achieve the highest performance and three datasets where they achieved comparable results to an existing model with varying levels of label noise. For a comprehensive analysis, we introduced label noise at various levels, including $5\%$, $10\%$, $15\%$, and $20\%$, intentionally corrupting the labels of these datasets. Table S.I displays the accuracies of all models for the selected datasets with $5\%$, $10\%$, $15\%$, and $20\%$ noise. Consistently, the proposed IFUTSVM-ID model demonstrates superior performance over baseline models, exhibiting higher ACC. Significantly, they maintain this leading performance despite the presence of noise. The average ACC of the proposed IFUTSVM-ID on the abalone9-18 dataset at various noise levels is $76.83\%$, surpassing the performance of the baseline models. On the checkerboard_Data dataset and ecoli-0-1_vs_5, the average ACC of the proposed model at different noise levels is $84.18$ and $78.47\%$, respectively. On the aus dataset, the proposed model achieves an average ACC of $85.05\%$, which is slightly lower than the average ACC of the IF-USVM model. However, the ACC of the proposed IFUTSVM-ID model reaches $86.47\%$ at a $0\%$ level of label noise. At each noise level, IFUTSVM-ID emerges as the top performer, with an overall average ACC of $79.79\%$. By subjecting the model to rigorous conditions, we aim to showcase the exceptional performance and superiority of the proposed IFUTSVM-ID model, especially in challenging scenarios. The above findings underscore the significance of the proposed IFUTSVM-ID model as resilient solutions, capable of performing well in demanding conditions characterized by noise and impurities.
Table S.I: Performance comparison of the proposed IFUTSVM-ID with the baseline models on KEEL datasets with label noise.
$ACC(\%)$ $ACC(\%)$ $ACC(\%)$ $ACC(\%)$ $ACC(\%)$ $ACC(\%)$
abalone9-18 $5\%$ $76.8$ $64.84$ $75.39$ $75.89$ $75.89$ $77.56$
$10\%$ $72.69$ $64.84$ $59.36$ $74.89$ $75.26$ $75.97$
$15\%$ $79.5$ $70.32$ $75.39$ $75.89$ $75.43$ $73.01$
$20\%$ $76.35$ $77.17$ $79$ $79.37$ $79.04$ $80.78$
Average ACC $76.34$ $69.29$ $72.29$ $76.51$ $76.41$ $76.83$
aus $5\%$ $84.37$ $85.51$ $85.99$ $85.51$ $85.47$ $86.98$
$10\%$ $85.41$ $85.51$ $85.99$ $69.08$ $84.37$ $86.47$
$15\%$ $83.92$ $67.15$ $85.51$ $65.7$ $84.41$ $84.54$
$20\%$ $85.44$ $59.42$ $85.51$ $82.61$ $82.61$ $82.19$
Average ACC $84.79$ $74.4$ $85.75$ $75.73$ $84.22$ $85.05$
brwisconsin $5\%$ $77.55$ $65.69$ $78.04$ $77.55$ $77.55$ $75.69$
$10\%$ $76.57$ $65.69$ $77.06$ $75.1$ $76.57$ $74$
$15\%$ $72.65$ $79.71$ $75.59$ $74.12$ $72.16$ $73.63$
$20\%$ $74.12$ $76.57$ $76.57$ $75.1$ $74.61$ $74.31$
Average ACC $75.22$ $71.92$ $76.82$ $75.47$ $75.22$ $74.41$
checkerboard_Data $5\%$ $82.37$ $81.51$ $81.99$ $82.51$ $85.47$ $85.51$
$10\%$ $85.41$ $85.51$ $85.99$ $69.08$ $86.37$ $86.47$
$15\%$ $82.92$ $67.15$ $80.51$ $65.7$ $81.41$ $84.54$
$20\%$ $80.44$ $59.42$ $80.51$ $79.61$ $80.61$ $80.19$
Average ACC $82.79$ $73.4$ $82.25$ $74.23$ $83.47$ $84.18$
ecoli-0-1_vs_5 $5\%$ $73.06$ $79.17$ $78.89$ $73.06$ $75.83$ $79.17$
$10\%$ $74.44$ $79.17$ $75.99$ $73.06$ $75.83$ $77.78$
$15\%$ $73.06$ $78.89$ $68.06$ $73.06$ $74.44$ $72.22$
$20\%$ $88.89$ $87.22$ $87.22$ $83.06$ $84.44$ $84.72$
Average ACC $77.36$ $81.11$ $77.54$ $75.56$ $77.64$ $78.47$
Overall Average ACC $79.3$ $74.02$ $78.93$ $75.5$ $79.39$ $79.79$
Table S.II: Comparison of the performance of the proposed IFUTSVM-ID model with the baseline models based on classification ACC on KEEL datasets.